### v1

# Policy Learning with a Language Bottleneck

Abstract
Modern AI systems such as self-driving cars and game-playing agents achieve superhuman performance, but often lack human-like features such as generalization, interpretability and human inter-operability. Inspired by the rich interactions between language and decision-making in humans, we introduce Policy Learning with a Language Bottleneck (PLLB), a framework enabling AI agents to generate linguistic rules that capture the strategies underlying their most rewarding behaviors. PLLB alternates between a rule generation step guided by language models, and an update step where agents learn new policies guided by rules. In a two-player communication game, a maze solving task, and two image reconstruction tasks, we show that PLLB agents are not only able to learn more interpretable and generalizable behaviors, but can also share the learned rules with human users, enabling more effective human-AI coordination. [^1] We provide source code for our experiments at https://github.com/meghabyte/bottleneck.
1 Introduction
As AI systems play an increasingly central role in automation, there is a growing need for them to better model and emulate human behavior. Such systems must be interpretable, and both act and generalize in human-predictable ways so that users can effectively interact with them. Unfortunately, many of today’s AI systems do not meet these standards. Self-driving cars or game-playing agents may achieve super-human performance but lack interpretability (McIlroy-Young et al., 2020) and often act unpredictably, especially outside their training distribution (Wang et al., 2023).
Figure 1: Policy Learning with a Language Bottleneck (PLLB) alternates between two steps: 1) gen_rule (bottom) generates a linguistic rule $\mathcal{L}_{i}$ explaining the agent’s best behaviors by prompting a language model with contrastive (positive and negative) episodes (middle); 2) update (top) learns a new policy conditioned on that rule. PLLB facilitates stronger human-AI collaboration, while constraining policies to be more interpretable and generalize better.
Humans acquire most of their skills and knowledge from others, often using language—via instructions, advice, or explanations that improve their decision-making capabilities (Carruthers and Boucher, 1998; Mesoudi and Thornton, 2018). Language acts as a communicative medium, enabling us to teach, learn from, and coordinate with others to solve complex problems. It also supports other cognitive functions: even when not used for communication, it allows us to represent abstract concepts (Hesse, 1988; Lakoff and Johnson, 2008), and plan (Vygotsky, 1965; Clark, 1998); it guides our attention (Waxman, 1994; Yoshida and Smith, 2003), and prompts relational thinking (Gentner and Loewenstein, 2002).
Consider a driver learning to navigate novel social conventions (e.g., triangle-shaped stop signs). While adapting to the environment, they might verbalize strategies to themselves to avoid future mistakes (e.g.,  If the sign is triangular, I should stop, a cognitive use), or transmit this convention to others (e.g.,  telling a friend In Japan, stop signs are triangles, a communicative use). Representing learned information in language helps humans solve problems and transmit knowledge by effectively capturing abstract problem structures that facilitate learning and generalization (Boutonnet and Lupyan, 2015; Chopra et al., 2019; Tessler et al., 2021).
The recent literature contains many examples of AI systems leveraging language-based representations or linguistic feedback, but these often rely on external supervision: humans in the loop or hard-coded feedback functions (see related work and reviews in Luketina et al., 2019; Colas et al., 2022). We argue that more human-like AI systems should not only use language-based supervision but also generate their own language-based feedback to leverage both the communicative and cognitive functions of language.
We introduce Policy Learning with a Language Bottleneck (PLLB), a framework providing artificial embodied agents the ability to generate linguistic rules that capture the strategies underlying their most rewarding behaviors. As shown in Figure 1, PLLB alternates between a gen_rule step that explains the agent’s most rewarding behaviors by prompting a language model (LM) with contrastive episodes, and an update step that learns a new policy conditioned on these rules. Unlike past work that solely leverages LMs for modeling agent behavior and multi-step reasoning (Park et al., 2023; Wei et al., 2023), PLLB is applicable even when aspects of the target policy cannot be expressed with language. The rules learned with PLLB can not only be shared with users to increase the agents’ interpretability and collaborative capabilities (communicative use), but also used internally by the agents to constrain their exploration, leading to more human-like and generalizable behaviors (cognitive use).
Policy Learning with a Language Bottleneck helps agents learn more human-like policies, which we examine across four diverse tasks. They perform better: in two image reconstruction tasks, PLLB agents generate instructions increasing the listeners’ performance compared to non-linguistic baselines (Section 7). They generalize better: in a maze task, agents uncover the abstract problem structure and generalize better to similar mazes (Section 6). They are more interpretable: in a coordination task, agents converge on the most human-interpretable policy when several optimal ones exist (Section 5). They are more inter-operable: in maze and image reconstruction tasks, humans achieve achieve better rewards when interacting with PLLB agents than agents trained without a bottleneck (Sections 6 and 7).
2 Background & Related Work
PLLB is inspired by the coupling of language and decision-making in both humans (Carruthers and Boucher, 1998) and machines (Colas et al., 2022).
Language for communication. Language facilitates cooperation and coordination between humans and machines via instructions (Hermann et al., 2017; Chevalier-Boisvert et al., 2019), advice (Watkins et al., 2021), explanations (Zhong et al., 2020; Lampinen et al., 2022), or the formation of conventions (Hawkins et al., 2020; Hu and Sadigh, 2023). Such communicative functions increase the fidelity and breadth of cultural transmission — a process of social learning that underlies human ecological success (Mesoudi and Thornton, 2018). PLLB agents not only learn from language, but also generate their own to be shared with others.
Language for cognition. Language also augments a learner’s cognitive abilities. Language-augmented RL agents represent more abstract goals (Jiang et al., 2019), generalize better (Hill et al., 2020; Colas et al., 2020; Wong et al., 2021), explore more efficiently (Colas et al., 2020; Tam et al., 2022; Klissarov et al., 2023) and can decompose complex goals into simpler ones (Chen et al., 2021; Ahn et al., 2022; Hu et al., 2022; Sharma et al., 2021; Hu and Clune, 2023). Our work extends these benefits to agents that learn from self-generated linguistic feedback.
Inner speech in humans. Generating linguistic rules for oneself is a form of inner speech. In the Vygotskian tradition, inner speech is seen as the internalization of the social speech generated by caretakers to help children solve problems (Vygotsky, 1965; Luria, 1959). As a result, it is thought to support our capacities for complex, long-term behaviors (Vygotsky, 1965; Luria, 1959; Hermer-Vazquez et al., 2001; Spelke, 2003).
“Inner speech” in machines. AI agents endowed with forms of inner speech (explanations, descriptions or subgoals) have been found to perform and generalize better than agents trained with purely neural representations (Wong et al., 2021; Lampinen et al., 2022; Roy et al., 2022; Hu and Clune, 2023). Unlike these approaches, PLLB approach generates language in an unsupervised way and maximizes downstream performance as well as interpretability and inter-operability.
Multi-step reasoning. Recent work has proposed to guide LMs’ reasoning by prompting them to step through sequences of “thoughts” (Wei et al., 2023; Yao et al., 2022; Li et al., 2023b; Shinn et al., 2024). Similarly, our approach uses a language bottleneck to concisely express intermediate information useful for later behavior (here policy learning). Unlike LM reasoning approaches, our generated language is used both for increasing interpretability and inter-operability of initially uninterpretable AI systems (low level policies) and for enchancing the performance and generalization capacities of an RL agent.
3 The Language Bottleneck
Figure 2: PLLB can be applied to a diversity of domains. In all domains, it iterates between gen_rule, a function that prompts an LM to extract a linguistic rule ($\mathcal{L}$, blue) by contrasting high and low reward episodes from past experience (green and red boxes), and update, a function for updating an agent’s policy through interaction with the environment conditioned on $\mathcal{L}$. PLLB can be applied to multi-step decision making tasks and visual tasks alike with minimal implementation variations (replacing the LM by a VLM in gen_rule, or replacing policy regularization by simple instruction conditioning in update. Exact prompts are in Appendix Section A.3.
PLLB builds on the standard RL framework to train agents to solve decision-making tasks in human-like ways. We formalize decision-making tasks (e.g.,  solving a maze) as Markov decision processes $(S,A,f,R,T)$ with reward function $R:S\times A\rightarrow\mathbb{R}$ (e.g.,  solving speed) over states $S$ (e.g.,  cell coordinates) and actions $A$ (e.g.,  directions), finite time horizon $T$, and a deterministic transition function $f:S\times A\rightarrow S$ that maps state–action pairs $(s,a)$ to next states $s$ (e.g.,  moving between cells). Standard RL then involves training a policy $\pi$ to maximize expected reward (e.g.,  by learning to solve the maze faster). This is usually done by alternating between two steps: (1) collecting data with the current policy: $D\leftarrow\pi_{i}$ and (2) updating the policy using the data: $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D)$, where update is an RL algorithm (Sutton and Barto, 2018). This procedure often fails to yield interpretable, inter-operable, or generalizable behaviors.
Our key idea is to introduce a language bottleneck between data collection and policy update. We extract linguistic rules explaining past rewarding behaviors and then use them to regularize the policy’s behavior in the next learning iteration (e.g., I should go right in blue cells, see Figure 1). Our experiments will show that this improves the interpretability, interoperability (communicative use), performance and generalization (cognitive use) of our agents over standard RL baselines in a variety of tasks (see Figure 2 and Sections 5 to 7). The resulting algorithm thus alternates between three steps: (1) data collection, $D_{i}\leftarrow\pi_{i}$; (2) language bottleneck generation, $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D_{i})$ and (3) policy updating, $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$, see Figure 1. In PLLB, the first step is implemented just as in ordinary policy learning approaches; below describe the last two steps in more detail.
3.1 Rule Generation (gen_rule)
Using all the experience $D_{i}$ collected by the policy $\pi_{i}$ in the current iteration, the gen_rule step aims at inferring an abstract rule $\mathcal{L}_{i}$ that best explains the agents’ successful behaviors: $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D)$. This is done by prompting an LM with contrastive episodes from $D_{i}$ (top-$N$ highest vs. top-$N$ lowest total rewards) and asking it to provide the rule that should be followed to obtain high rewards (see first row of Figure 2 and Appendix Section A.3 for full prompts). We found this contrastive approach, inspired by Zhong et al. (2023) and Dunlap et al. (2023), to outperform simple summarization of high-reward samples.
3.2 Rule-Guided Policy Update (update)
Given a rule $\mathcal{L}_{i}$, the update step produces a new policy $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$ that is better aligned with $\mathcal{L}_{i}$. One existing approach to solving this problem is the InstructRL approach of Hu and Sadigh (2023), which regularizes the learned policy with another policy induced by the linguistic rule $\pi_{\mathcal{L}}$. For instance, in the maze example shown in Figure 1, if the rule is I should go right on every blue cell, the induced policy $\pi_{\mathcal{L}}$ should assign probability 1 to the right action in blue cells, but equal probabilities to all actions in every other situations. In the Q-learning algorithm, this approach simply adds a regularizing term (orange) to the standard Q-learning update rule  [^2] This regularizing term can also be applied in the function approximation setting (Mnih et al., 2013).:
$Q^{\theta}(s_{t},a_{t})\leftarrow r_{\text{t+1}}+\gamma Q^{\theta}(s_{t+1},a_{t+1}),\textrm{ where}$
$a_{t+1}=\arg\max_{a}[Q(s_{\text{t+1}},a){\color[rgb]{1,.5,0}+\lambda\log\pi_{\mathcal{L}}(a\mid s_{t})}].$
Here $\gamma$ is a discount factor and $\lambda$ controls the strength of the rule-induced regularization and could be made time-dependent with a pre-defined or learned schedule (e.g., stronger regularization at the beginning of learning).
But how do we induce $\pi_{\mathcal{L}}$ from $\mathcal{L}$? Since $\mathcal{L}$ is expressed in natural language, we emulate it with an LM. We prompt it with both the rule and the current state of the agent $s_{t}$ and instruct it to generate the next action to obtain a probability distribution over admissible actions (e.g directions). This can be done by normalizing the probability of generating each of the action tokens, or by counting the frequency of these tokens over multiple LM generations. Running Q-learning updates with the regularization term from experience data gives us a new Q-table $Q^{\theta}_{i+1}$ from which we can derive the new policy $\pi_{i+1}$ by taking actions with maximum expected value in every state $\pi_{i+1}(s_{t})=\text{max}_{a}Q^{\theta}_{t+1}(s_{t},a)$.
As an alternative, in certain domains, the policy can be directly implemented by an LM or VLM. In that case, simply conditioning the policy on the rule by adding it to the prompt $\pi_{\mathcal{L}}$ might be enough to steer the agent’s behavior to become more human-like: $\pi_{i+1}\leftarrow\pi_{\mathcal{L}_{i}}$. Given a new policy $\pi_{i+1}$, we can now close the loop and generate a new rule $\mathcal{L}_{i+1}$. We implement PLLB with policies from RL agent exploration (Sections 5 and 6) and LM/VLMs (Section 7) in our experiments.
4 Experiment Set-Up
We run experiments on a set of four diverse tasks to showcase the capacity of PLLB to train more human-like agents. Our domains include a simpler two-player communication game called SelectSay (Section 5), a Maze solving task (Section 6), and two collaborative image reconstruction tasks with either synthetic (Builder) or natural (Birds) images (Section 7). For all tasks, the exact prompts used in gen_rule and update and examples of rules generated across iterations can be found in Appendix A.3 and B.1 respectively.
To evaluate the cognitive function of PLLB, we first evaluate the policies learned by artificial agents in each environment based on performance metrics such as task reward, similarity to known human-like policies, or generalization. For all tasks, we compare our Bottleneck approach with baseline non-linguistic methods (e.g., standard RL, base LM) and methods leveraging invalid linguistic rules (e.g., obtained by labeling high reward examples as low reward) in gen_rule (Adversarial). These baselines allow us to study the impact of generating linguistic rules aligned with the underlying environment structure.
To evaluate the communicative function of PLLB, we conduct a variety of human subject studies showing that generated rules $\mathcal{L}$ can help improve human task performance on a variety of metrics (e.g., reward, solving time). Human subject study participants were either recruited in-person (for Maze) or on the Prolific crowdsourcing platform [^3] Prolific, http://prolific.com. (Builder, Birds), compensated at a rate of US\$12/hour. All studies were approved by our institution’s IRB.
Figure 3: Overview of SelectSay game, reproduced from Hu and Sadigh (2023).
5 SelectSay
As a proof-of-concept, we first consider SelectSay, a simple collaborative game introduced in Hu and Sadigh (2023). Here, a speaker can see a hidden set of five balls including two blue ones and must help a listener find the two blue balls by communicating with them only via numbers, see Figure 3. In principle, the speaker and listener could converge on any bijective convention between messages and balls, and this is what we find with RL agents (multiple optimal policies). Humans, on the other hand, would prefer the intuitive mapping 1 $\rightarrow$ 1, 2 $\rightarrow$ 2, etc.  that we call the human-interpretable policy. Hu and Sadigh (2023) showed that regularizing the listener’s policy with the instruction I should select the same number as my partner, successfully guided the two RL agents to the human-interpretable policy. We investigate whether PLLB could generate a similar rule and achieve a similar interpretability on its own.
Figure 4: SelectSay. a) Bottleneck agents learn as fast as TabularQ and InstructRL and faster than agents using adversarial rules Adversarial. b) Unlike TabularQ, they learn human-interpretable policies; this without relying on an external instruction like InstructRL. c) When faced with speakers enforcing a non-human-interpretable policy, Bottleneck converges faster than InstructRL.
Task Overview. Like in Hu and Sadigh (2023), the speaker and listener are RL agents trained with Q-Learning that receive positive rewards when the listener collects blue balls, negative rewards otherwise. After training both agents for 200 episodes, we prompt llama-2-70b-chat [^4] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf with the task description and a contrastive set of high- and low-reward episodes to generate a rule $\mathcal{L}$, e.g., I should choose action 4 whenever the observation is 2, 3, 4, or 5. (gen_rule, see Figure 2).
We generate new rules every 500 training episodes. In between, we regularize the listener’s policy with a another policy $\pi_{\mathcal{L}}$ induced from the current rule $\mathcal{L}$. This is done by applying a softmax on the LM’s logits (see method in Section 3) across all possible actions and obtaining a distribution. In the following experiments, we compare our Bottleneck method, with the Adversarial version (corrupted rule), a TabularQ baseline, as well as an InstructRL upper bound using the same predefined instruction as (Hu and Sadigh, 2023).
5.1 PLLB helps learn human-interpretable policies
We first compare the four methods along two dimensions: their reward and their interpretability measured as the proportional (0 to 1) similarity between the listener’s policy and the optimal human-interpretable policy (the intuitive message-ball described above). All methods but Adversarial quickly solve the task (Figure 4a), showing that corrupted rules hinder learning. Bottleneck and InstructRL both converge on the human-interpretable policy thanks to their linguistic rules while TabularQ — deprived of such rule — does not (Figure 4b). Interestingly, we note that Adversarial policies are more interpretable than TabularQ ones, suggesting that even corrupted language may provide inductive biases towards more human-like behavior. Qualitatively, we observe that generated rules converge towards describing the human-interpretable policy (e.g.,  I should follow the strategy of choosing the same action as Agent 1, see examples in Appendix Section B.1).
5.2 PLLB can learn counter-intuitive policies
A potential confound in the preceding experiment is the possibility that PLLB helped converge to the human-interpretable policy because this is the only behavior an LM is able to describe. To test this hypothesis we fixed the speaker to use a counter-intuitive policy using a random mapping between balls and messages (e.g.,  1 $\rightarrow$ 3, 5 $\rightarrow$ 2, etc.). Here, a listener following the human-interpretable policy would fail. In Figure 4c we observe that while InstructRL and Adversarial, both regularized by misaligned rules (the human-interpretable one for InstructRL, another random permutation for Adversarial), are able to eventually adapt to the fixed speaker, they converge at a slower rate than both TabularQ (no rule) and Bottleneck. Overall, Bottleneck is the only algorithm able to flexibly adapt to a fixed speaker policy while converging on the human-like policy when the speaker is also learning.
6 Maze
Figure 5: Results in Maze. a) Bottleneck agents learn as fast as the non-linguistic Baseline agents, but faster than Adversarial and LinearQ agents. b) When faced with a new maze with similar structure, Bottleneck agents learn faster than TabularQ (which does not perceive color) and LinearQ(which does, but learns slower). c) When faced with a maze of a different structure, Bottleneck agents adapt swiftly while LinearQ cannot recover. We do not evaluate Adversarial for the Generalization or Adaptation experiments due to its poor performance in the standard setting.
In the Maze domain, agents must navigate a maze to find the goal using four directional actions (N/S/E/W). We study mazes in which environment cues provide hints about the optimal path. This setup helps us study the impact of the language bottleneck: can PLLB agents infer the underlying structure and use it to generalize across mazes (cognitive use)? Can they transmit this knowledge to humans and help them perform better in novel settings (communicative use)?
Task Overview. We randomly generate different 7x7 mazes with the gym-maze code base. [^5] https://github.com/MattChanTK/gym-maze Agents are rewarded when they achieve the goal, by the inverse of the number of steps they needed to reach it as an incentive to learn the optimal solving path. We introduce additional structure to all mazes by coloring cells for which the optimal action is south (red) or north then east (blue) with 50% probability (or they are left blank). Therefore, agents that leverage color information may better generalize to new mazes where the coloring structure is preserved, but the optimal action sequence differs.
The first iteration of PLLB correspond to training a standard RL algorithm to obtain $\pi_{1}$, in our case a tabular Q-learning agent (TabularQ), since there are no rules yet. After the agent observes at least two solved mazes, we run gen_rule by prompting a llama-2-70b-chat LM with contrastive episodes and a task description. This gives us a linguistic rule $\mathcal{L}_{1}$ (e.g., Upon observing RED, take SOUTH) that we can use to update the policy. We first induce the regularizing policy $\pi_{\mathcal{L}_{1}}$ with rule $\mathcal{L}_{1}$ by obtaining a probability distribution over the 4 actions from the LM, and then run the RL algorithm for 5 episodes to obtain the new policy $\pi_{2}$ (procedure described in Section 3). We repeat these steps every 5 episodes of interactions with the environment.
Baselines include the base tabular Q-learning algorithm without language bottleneck (TabularQ) and a variant of Bottleneck generating rules from reward-randomized episode samples (Adversarial). We additionally evaluate LinearQ, an agent learning a linear model Q-function with an additional feature for cell color. We train LinearQ with a batch update of size 10 and learning rate 0.001, after performing a hyperparameter sweep, and note that we found it generally took significantly longer to converge than the other methods.
6.1 PLLB learns more generalizable policies
Figure 5a shows that learning a valid rule (Bottleneck) does not increase learning speed over TabularQ (already quite fast), but using a corrupted rule (Adversarial) or learning from linear features (LinearQ) slows down learning. We next replace the maze by another 7x7 maze sharing the same underlying color semantics, but with a different optimal action sequences. For fair comparison, we start with the fully converged policy for each method (requiring 100 episodes for LinearQ). Figure 5b shows that Bottleneck leverages the learned rule and adapts to the new maze more effectively than both TabularQ and a fully converged LinearQ agent. While TabularQ cannot generalize because it does not perceive colors, LinearQ does generalize faster at first but converges slower.
We find that the generated rules indeed improve over time to better capture the underlying structure of the maze (e.g., if I observe BLUE, then take the NORTH action; see other examples in Appendix  B.1). Across all trials, $100\%$ of the final rules mention the red $\rightarrow$ south rule and $60\%$ uncovered the more complex blue $\rightarrow$ north-then-east rule.
6.2 PLLB learns adaptable policies
In Figure 5c, we reproduce this same experiment on a maze with a different underlying structure (red now indicates west while blue indicates east then south). Although the rule Bottleneck learned does not apply anymore, it can still adapt faster than the baselines. The language bottleneck can indeed quickly capture the new structure of the task and regularize learning to steer adaptation. Here LinearQ seems to have overfit to the first maze and struggled to adapt. We find that all trials end with rules capturing the new mapping (100% red $\rightarrow$ west, 50% blue $\rightarrow$ east-then-south). Overall, these results show that the language bottleneck supports human-like cognitive functions by finding a tradeoff between the efficiency of TabularQ and the generalization capabilities LinearQ while remaining more adaptable.
6.3 PLLB is more interpretable and inter-operable
Figure 6: Participants provided rules from PLLB solve new mazes faster than those given either visual or no aid (control), left. They also self-reported linguistic feedback as more useful than visual (right).
Can generated rules be useful to humans as well? We asked 50 university students to solve a 7x7 maze in the fewest steps possible using arrow keys. They could only observe the cells they had already visited and the walls they had already bumped into (see Figure 12 in Appendix). We split them into three groups: a Control group receiving no assistance, and two others receiving information about a similar but different maze sharing the same underlying color semantics. The Visual group is shown a visual representation of the optimal policy (arrows in each cell, see Figure 11 in Appendix), while the Bottleneck group is provided a randomly sampled language rule $\mathcal{L}$ learned by PLLB. This set-up lets us evaluate how generalizable the two different aids are to a new maze, as well as how quickly participants can account for incorrect information depending on modality.
Participants using PLLB rules solve the new maze with fewer steps than others and find this aid more useful than the visual one in average (statistically significant at level 0.05 with two-sided Mann-Whitney test after Bonferroni correction, see Figure 6). Participants indeed found it harder to extract the relevant information from the visual aid, which contained a great deal of non-transferable information (optimal actions in all non-colored cells). PLLB rules focus on the useful and transferable information learned by the previous agent, which is much more readily usable by humans.
Overall, the results in Maze demonstrate the ability of PLLB to train agents that are more generalizable and adaptable (cognitive use) as well as more interpretable by, and inter-operable with humans (communicative use).
7 Collaborative Image Reconstruction
PLLB is not restricted to the training of RL agents. This section introduces two collaborative image reconstruction tasks inspired by the collaborative assembly task of McCarthy (2020). Builder and Birds both consider two agents: a speaker, who can see a hidden target image, and a listener, who must accurately reconstruct the target image based on a description from the speaker. Both agents aim to converge to a description style leading to high reconstruction accuracy. We consider both synthetic images built from sequences of discrete actions by the listener (Builder), and natural images of birds that the listener reconstructs using a text-to-image generation model (Birds). Here, PLLB combines with state-of-the-art multi-modal models to let agents perceive semantic visual features.
7.1 Task Overview
Dataset.
In Builder, we construct a dataset of synthetic images with a variable number of shapes (triangle, square, circle) in different colors (magenta, blue, red, green) and locations on a 2D grid. Each image is created by a sequence of discrete actions, each representing a particular combination of shape, color, and area of the grid (e.g., [ACT12 ACT2 ACT4], where ACT2 adds a magenta circle in the lower right area of the grid). In Birds, we construct a dataset of natural bird images by selecting 5 images for each of 10 bird species from the CUB-200-2011 dataset provided by Wah et al. (2011) (see Appendix Section A.2 for species list).
Reward.
In Builder, because target and reconstructed images are completely defined by the action sequences that generate them, we measure task success with the Levenshtein similarity between the (sorted) corresponding action sequences. However, the listener in the Birds environment outputs images using a text-to-image model that exhibit a variety of properties. We therefore introduce reward functions corresponding to three different properties for Birds: Color, Background, and Species. To evaluate a listener’s reconstructed image $x_{r}$ with respect to a particular target image $x_{t}$ and reward function, we create a contrast set $C$ by selecting the three images in CUB-200-2011 dataset that have a different value for the target reward property (e.g., color), but are most similar with respect to values for all other annotated properties. We then define reward $r$ as $r=\sum_{c\in C}d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{c}))-d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{t}))$, where CLIP refers to the CLIP image embedding (Radford et al., 2021). For both Builder and Birds tasks, reward is reported over a separate held-out validation set.
Models.
As open-source VLMs struggle with describing synthetic images, we use the llama-2-70b-chat LM as our speaker for Builder, representing images in raw text that list all shapes’ type, color, and exact coordinates individually in sequence. On the other hand, we use the open source Llava VLM as our speaker for Birds (Liu et al., 2023). We prompt both speaker models using the target image and a general task description, as well generated rules after the first iteration.
We implement the listener for Builder with BART, a neural sequence-to-sequence model pre-trained on English text by Lewis et al. (2019), which we fine-tune at each episode on a training set of (description, action sequence) pairs using descriptions provided by the speaker. Meanwhile, we use the Stable Diffusion text-to-image diffusion model as the listener for Birds.
For both tasks, we consider two settings: From Scratch, where the listener is initialized using the original pre-trained model weights at every episode, and Continual Training, where the listener is continuously updated over time using descriptions by the speaker. We therefore split our datasets into separate held-out sets for training, early stopping, selecting samples for gen_rule, and evaluation.
Implementation.
We first generate base image descriptions for both tasks by providing the speaker a prompt containing a general task description and target image, but without any rule $\mathcal{L}$. For the Builder task, an original image description is: At least five dots in total - four red and one green - scattered across the canvas, with two sets of matching locations (x = 0.49 / y = 0.24 and x = 0.5 / y = 0.24), another set at x = 0.52 / y = 0.3, and finally, one dot located near x = 0.76 / y = 0.71. This description is overly complex, mixes different coordinate formats, and difficult for the listener to use to reconstruct the target image, leading to low reward. Meanwhile, an example original image description for the Bird task is: “a bird on a branch”. While simple, this description does not contain sufficient information about the Color, Background, and Species, causing the listener to generate an image reconstruction with low reward.
We next evaluate these base descriptions using each task’s respective reward functions, and select the (image, description) examples with the 5 highest and lowest rewards. Using the prompts shown in Section A.3, we implement gen_rule using the llama-2-70b-chat and llava models for Builder and Birds, respectively. For Birds, we observe that output rules $\mathcal{L}$ are specific to the reward function: an example rule we get for the Species reward is: “Identify the bird’s species if possible, and include any distinctive characteristics that set it apart from other birds.”, while an example rule we get for the Background reward is: “Describe the setting or background of the image, such as the presence of snow, water, or other elements”.
We implement update for both tasks by simply appending the output rule $\mathcal{L}$ to the original prompt to the speaker, as shown in Figure 2, using the speaker’s output as the new description for a given target image. We repeat this cycle of eliciting rule $\mathcal{L}$ , appending $\mathcal{L}$ to the same prompt used to generate the base instructions in order to re-label our data with new instructions, and training and evaluating the listener model for a total of 5 iterations for Builder and 3 iterations for Birds. We qualitatively show this leads to improved image reconstruction from the listener in Figure 7, and next discuss our experimental results.
Figure 7: Both the builder and birds tasks consist of speaker and listener agents. At test time, the speaker needs to provide a language description to the listener that helps them recreate the image accurately. For both tasks, PLLB helps improve listener accuracy over time.
Figure 8: In Builder, performance of the Bottleneck listener improves upon Baseline descriptions over time with PLLB for both From Scratch and Continual Learning training settings, unlike an (Adversarial) listener that inaccurately represents rewards.
7.2 PLLB helps speakers provide more usable instructions
For both tasks, listeners following rules generated by PLLB (Bottleneck) outperform listeners trained from uninformed (Baseline) or misinformed (Adversarial) speakers (Figures 8 and 10, top row). This holds true for all three different reward functions in Birds, as well as both training settings, with Continual Learning learning enabling Bottleneck to have an even stronger improvement over Baseline image descriptions by leveraging learned task experience. Continual Learning does not significantly help the Adversarial method, emphasizing that the linguistic rules $\mathcal{L}$ in Bottleneck do capture succesfull task strategies. However, Continual Learning is not required for Bottleneck to outperform Baseline image descriptions. The success of the From Scratch training setting can be interpreted as the speaker generating abstract rules to guide its own learning, leading to improved performance and showing evidence for a cognitive use of language.
Figure 9: In Builder, study participants reconstructed target images with higher accuracy (left), and less time (right), when provided instructions that were generated using Bottleneck vs. original Baseline descriptions.
Finally, we observe that the generated rules $\mathcal{L}$ encourage the speaker to reduce vague and general language, and describe properties relevant to the corresponding underlying reward function in Birds, although this might take several cycles to fully converge (see example rules in Appendix Section B.1). Furthermore, rules $\mathcal{L}$ become more complex over time, and across different trials we observe path dependency where different trials converge to different rules (e.g., $(x,y)$ vs. $x=0.xx,y=0.yy$ for coordinates in Builder). Overall, these rules help guide the speaker and listener agents in both tasks to iteratively improve image reconstruction over time, which can also be seen qualitatively in Figure 7.
7.3 PLLB can collaborate with human listeners
Do human listeners also benefit from image descriptions provided by speakers following Bottleneck rules? We next conduct a human subject study for the Builder task, where we replace the listener with study participants at the end of one iteration of PLLB, and evaluate how accurately and quickly they can reconstruct target images from a held-out test set. We recruit 20 crowdworkers on Prolific to reconstruct 5 images using descriptions from a speaker following a rule $\mathcal{L}$ generated with (Bottleneck), and 5 images using the original (Baseline) descriptions. We provided participants an interface that included a drawing canvas and buttons controlling shape types and colors (Section A.1), allowing us to use our reward function of Levenshtein similarity between user actions and the underlying action sequences for a target image.
Figure 10: For the Birds task, rules generated after 3 iterations of Language Bottleneck help automated (top) and human (bottom) speakers generate better descriptions for naturalistic image reconstruction for three different underlying reward functions (columns). Continual learning (dark teal, red in top graphs) further improves reward over listener agents learning from scratch (dark teal, red in top graphs), but not sufficiently enough to help Adversarial outperform baseline descriptions, showing that Bottleneck rules describe rewarding behavior.
Participants using Bottleneck descriptions built target images faster and more accurately than participants using control descriptions (Bonferroni-corrected Wilcoxon signed-rank test, $p<0.05$, Figure 9). 55% of them preferred Bottleneck descriptions when asked in a post-study survey, describing them as “more direct and less ambiguous”. 10% answered that both descriptions types were equally easy to follow while remaining participants preferred Baseline descriptions because they they gave more flexibility, showing a preference not captured in our reward function.
7.4 PLLB collaborates with human speakers
Can humans benefit from rules generated by PLLB? We asked 12 Prolific crowdworkers to act as speakers and provide descriptions for our dataset of Birds images. Half of the participants were provided rules corresponding to one of the three reward functions ($\mathcal{L}_{3}$) while the other half were not. Figure 10 (bottom) shows that listeners instructed by human speakers provided with PLLB rules outperform those instructed by uninformed human speakers (Baseline) (Bonferroni-corrected t-tests $p<0.05$ for species and colors).
The descriptions of workers not provided a rule are more diverse and less focused e.g., A barbed wire is an uncomfortable stop for a bird. With PLLB rules on the other hand (e.g., Describe the bird’s coloration accurately), they generate more specific descriptions: Red and black bird on barbed wire. Bright red chest, red at top of head, black wings and beak for the same target image (see other example rules in Appendix Table 4). PLLB agents can easily transmit what they learned from experience (the rule) to humans who can then benefit from that knowledge (cultural transmission, communicative use).
8 Limitations & Discussion
By inferring the rule explaining its best past behaviors, PLLB becomes more interpretable by and inter-operable with humans. When it uses a rule to condition its policy updates, it steers its own behavior to become more interpretable and generalizable. In our current implementation, rule generation remains limited: we require converting agent histories into representations the LM can handle (e.g., text or images), which may prevent the use of long-horizon sensorimotor trajectories (e.g., robotic systems). Rule generation can also fail when the LM overfits or abstracts too much (e.g., in Birds we found the rule: “Use descriptive language that conveys the mood, such as the serenity of a snowy day or the freedom of flight”). Advances in multi-modal models and better approaches for selecting contrastive examples for gen_rule might help mitigate these issues.
An interesting direction for future work is to apply PLLB to tasks with complex reward functions that include hard-to-articulate human preferences. In applications like image captioning for accessibility (Nie et al., 2020) and personalization of language models (Li et al., 2023a), linguistic rules might improve transparency and predictability even in cases where users find it difficult to describe their own goals.
Finally, we could extend PLLB to generate other kinds of linguistic information: not only abstract rules, but also explanations, goals, or learning strategies. Such information could then be shared with others (cultural transmission) or used by update functions in new ways: mapping generated language to state or action abstractions, goals, reward functions, or even learning strategies. This opens the door for more human-like AI agents leveraging the deep interactions between language and decision-making (Colas et al., 2022).
Acknowledgements
We would like to acknowledge support from NSF Awards #2006388, #2125511, #2238240 and #2212310, AFOSR YIP, ONR #N000142112298, and DARPA. Megha Srivastava was additionally supported by an IBM PhD Fellowship. Cédric Colas received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065949.
Appendix A Additional Methods
A.1 Maze
Figure 11: Visual aid provided to participants in the human subject study for the Maze task.
Figure 12: Interface used in the human subject study for the Maze task. Cells not visited by the participants are hidden in gray.
A.2 Dataset Details for Birds.
We consider images of the following bird species in the CUB-200-2011 dataset from Wah et al. (2011) for the Birds task: [Indigo Bunting, Cardinal, Yellow Breasted Chat, American Crow, Vermillion Flycatcher, California Gull, Blue Jay, Tropical Kingbird, White Pelican, Horned Puffin].
Figure 13: Interface used in the human subject study for the Builder task.
A.3 Full Prompts
We first provide the full prompts used for gen_rule. For space constraints, we do not include example samples, but note they follow the format shown in Figure 2.
SelectSay:
You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Possible ACTIONS an agent can take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION describes the ordered sequence of actions that AGENT 1 picks, and each ACTION describes the ACTION that AGENT 2 picks based on the given OBSERVATION. The examples are separated into HIGH REWARD and LOW REWARD examples.+[samples]+Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
Maze:
You will be given a list of example (OBSERVATION, ACTION) trajectories collected from an AGENT learning to solve a maze. Each trajectory receives a REWARD. Possibles OBSERVATIONS an agent see are: WHITE, RED, BLUE Possible ACTIONS an agent can take are: NORTH, SOUTH, EAST, WEST. The examples are separated into HIGH REWARD and LOW REWARD examples + [samples] + Output a language rule that best summarizes the strategy the AGENT should follow when picking a sequence of ACTIONS to solve the maze and receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
Builder:
There are two agents. The goal of Agent 1 is to provide instructions to Agent 2 that helps Agent 2 to successfully recreate the image. You will be given a list of (ORIGINAL, AGENT 1 INSTRUCTION, REWARD) values where ORIGINAL is the original description of an image, INSTRUCTION is the instruction provided by Agent 1 to Agent 2, and REWARD is the reward Agent 2 receives when trying to re-create the image (higher is better). The examples are separated into HIGH REWARD and LOW REWARD examples. + [samples]+
Based on the examples above, output a list of 2 RULES for Agent 1 to follow when generating INSTRUCTION in order to receive HIGH REWARD, instead of LOW REWARD. Write the rules after the prefix RULES:
Birds:
The top row of three images have the following HIGH REWARD descriptions:+high reward samples+The bottom row of three images have the following LOW REWARD descriptions:+low reward samples+Provide a rule I should follow in order to provide image descriptions with HIGH REWARD, not LOW REWARD. Provide the rule after the prefix RULE:"
We next provide the full prompts used in update for each task.
  1. 1.
    SelectSay: $[\mathcal{L}]$+Agent 1 selected [observation]. So I should select
  2. 2.
    Maze: You are an agent solving a maze following a provided RULE. You will be given a list of PREVIOUS ACTIONS and the CURRENT OBSERVATION. Follow the RULE to select your NEXT ACTION (East, West, South, North):
    RULE: + $[\mathcal{L}]$+ PREVIOUS ACTIONS: + $[\tau_{1...t-1}]$ + CURRENT OBSERVATION: + [observation] +
    What is the NEXT ACTION you should take? Output one of (East, West, South, North) after the prefix NEXT ACTION:.
  3. 3.
    Builder: You will be given a DESCRIPTION of an image. Your goal is to use this description to provide a short INSTRUCTION to help someone else, who cannot see the image, accurately re-construct it. You will also be given a list of RULES you must follow when providing the instruction.
    DESCRIPTION: + observation +
    RULES: $+\mathcal{L}$
    Please provide a short instruction following the prefix INSTRUCTION:
  4. 4.
    Birds: Provide a one-sentence description of this image, using the following $\mathcal{L}$
Appendix B Additional Results
B.1 Examples of Generated Rules for All Environments
Setting $\mathcal{L}_{1}$ $\mathcal{L}_{2}$ $\mathcal{L}_{5}$
Standard I should follow the strategy of choosing action 4 whenever the observation is 2, 3, 4, or 5. I should follow the strategy of choosing actions based on the observation provided by AGENT 1. If the observation is 1, I should choose action 1. If the observation is 2, I should choose action 3. If the observation is 3, I should choose action 3. If the observation is 4, I should choose action 4. If the observation is 5, I should choose action 5. I should follow the strategy of choosing the same action as AGENT 1 for observations 1, 2, 3, 4, and 5.
Fixed Speaker I should choose action 1 when observation is 1 or 2 or 4 or 5. I should choose action 2 when observation is 3. I should choose action 1 when observation is 1, 2, 3 or 5. I should choose action 2 when observation is 4. I should choose action 1 when observation is 1. I should choose action 4 when observation is 2. I should choose action 5 when observation is 3. I should choose action 2 when observation is 4. I should choose action 3 when observation is 5.
Table 1: Example $\mathcal{L}$ rules generated for the SelectSay environment, for the Standard setting (both Listener and Speaker agents are RL agents trained from random initialization) and a Fixed Speaker agent.
Size $\mathcal{L}_{1}$ $\mathcal{L}_{2}$ $\mathcal{L}_{5}$ $\mathcal{L}_{8}$
7x7 I should start by moving north, then continue moving north until I reach a wall, then turn east and move until I reach a wall, then turn south and move until I reach a wall, then turn west and move until I reach the end of the maze. I should prioritize moving North, East, and South over moving West, as the HIGH REWARD EXAMPLES show a higher frequency of these actions. 2. I should avoid moving West unnecessarily, as it is rarely seen in the HIGH REWARD EXAMPLES and often leads to LOW REWARD EXAMPLES I should take the EAST action when I see WHITE, and then repeat the EAST action until I see RED, then take the SOUTH action, and then repeat the EAST action until I see BLUE, then take the NORTH action, then take the EAST action until I see RED again, then take the SOUTH action, and finally take the WEST action. I should take the EAST action when I observe WHITE, and keep taking EAST actions until I observe RED, then take the SOUTH action. After taking the SOUTH action, I should take the EAST action again until I observe BLUE, then take the NORTH action. After taking the NORTH action, I should take the EAST action until I observe RED again, then take the SOUTH action.
Table 2: Example $\mathcal{L}$ rules generated for the Maze environment for two different maze sizes.
Setting $\mathcal{L}_{1}$ $\mathcal{L}_{2}$ $\mathcal{L}_{3}$
Re-Initialization 1. Be specific with location details: Agent 1 should provide detailed location information for each element in the image, such as x and y coordinates. 2. Use descriptive language for elements, such as "red dot" or "green triangle". 1. Use specific coordinates when instructing Agent 2 to draw shapes. 2. Use descriptive language to specify the color and shape of each element. For example, "a green triangle" instead of "a green thing". 1. Be specific with location coordinates: provide specific coordinates for the location of each shape, using the format x=0.XX, y=0.YY. 2. Use descriptive shape names: Instead of using generic terms like "dot" or "square," use more descriptive names that indicate the shape’s color and size, such as "green triangle" or "red square."
Continual Training 1. Be specific and detailed in your instructions. High reward examples have specific coordinates and shapes, while low reward examples have more general descriptions. 2. Use a consistent format for your instructions. High reward examples have a consistent format for listing coordinates and shapes, while low reward examples have a more free-form format. 1. Provide explicit coordinates for each element in the image, using the format (x, y). 2. Use specific colors when referring to elements in the image, such as "red", "green", or "blue". Avoid using vague terms like "colored" or "shaded". 1. Use a consistent format for describing shapes, such as always listing the x-coordinate first, followed by the y-coordinate. For example, instead of "one green square at the point x=0.53, y=0.24", use "one green square at (0.53, 0.24)". 2. Avoid using vague terms like "various shades of green". Instead, use specific colors, such as "green" or "blue". Additionally, use specific shapes, such as "square" or "triangle", rather than vague terms like "rectangle".
Table 3: Example $\mathcal{L}$ rules generated for the Builder environment.
Reward $\mathcal{L}_{1}$ $\mathcal{L}_{2}$ $\mathcal{L}_{3}$
color Describe the bird’s color, species, and any distinctive markings or patterns. Describe the bird’s coloration accurately. Describe the bird’s coloration accurately.
background Include details about the bird’s surroundings, such as the type of branch or post it is on, and any additional elements in the background. Include the bird’s action (perched, flying, standing) and its location (on a branch, railing, pole, etc.) Describe the bird’s action (flying, perching, standing) and the environment it is in (sky, tree, water).
species Describe the bird’s color, markings, and any distinctive features. Describe the subject’s unique features, such as coloration, beak shape, or other distinguishing characteristics. Include specific details about the bird’s appearance, such as the color of its feathers, beak, or eyes, and any distinctive markings or patterns.
Table 4: Example $\mathcal{L}$ rules generated for the Birds environment demonstrate reward-specificity over time.